Two marbles are drawn without replacement from a jar with 4 black and 3 white marbles. What is the probability that both marbles are white? The the second marble drawn is white, given that the first is black? The the first marble is black and the second marble is white? and at least one of the marbles is black?
Answer by stanbon(75887)

Two marbles are drawn without replacement from a jar with 4 black and 3 white marbles.
What is the probability that both marbles are white?
Ans: 3C2/7C2 = 3/21 = 1/7
---------------------------------
That the second marble drawn is white, given that the first is black?
Ans: P(w|b) = 3/6 = 1/2
------------------------------
That the first marble is black and the second marble is white?
Ans: P(b and w) = (4/7)(3/6) = 2/7
------------------------------
That at least one of the marbles is black?
Ans: P(condition) = 1 - P(2 white) = 1 - 3C2/7C2 = 1-[3/21] = 6/7
=======================
Cheers,
Stan H.
====================
-------------------